# In situ calibration of microwave attenuation and gain using a cryogenic on-chip attenuator

### v1

@@ Abstract
Figure 2: a) Sketch showing the frequency overlap between the thermal noise, the transmission band of the amplification chain, and the three output band-pass filters (1: 6.8-7.8 GHz, 2: 4.9-6.2 GHz; 3: 3.5-4.5 GHz) used to restrict the analog bandwidth in different power spectral density (PSD) measurements. For each filter, the signal tone to be calibrated, $\omega_{\rm sig}$, is swept in the rejection band of the filter (gray region), whereas noise is detected at $\omega_{\rm det}$ (black dot) in the transmission band (colored region). b) Power spectral densities measured at $\omega_{\rm det}=7$ GHz by Joule heating (purple) and a RF heating (blue) at $\omega_{\rm sig}=4$ GHz of the on-chip attenuator. The tone applied was rejected at room temperature with the output band-pass filter 1. The Joule and RF heating traces are separated by the attenuation $A$ of the input line. The tone heating was also acquired via the reference line without the on-chip attenuator (orange) to show the onset of heating of the bulkhead attenuators (diamond marker). Since their contribution is negligible (see main text), the best overlap (red) of the RF heating with the Joule heating giving the attenuation $A$ was fitted from $P_{\rm sig}=-22.5$ dBm (above the noise floor, dot marker) until an arbitrary power $P_{\rm sig}=-5$ dBm (cross marker). c) Attenuation profile obtained by comparing Joule and RF heating at different $\omega_{\rm sig}$. Three detection frequencies were considered and implemented with the set of output band-passes depicted in a). The red band in the background highlights the bandwidth of the amplification chain.
@@
where $N_{\rm sig}$ is the number of photons at frequency $\omega_{\rm sig}$ at the input of the amplification chain, SNR the signal to noise ratio and $P_{\rm input}=A_{\rm line}A_{\rm att}P_{\rm sig}$ the power at the input of the amplification chain. The tone power $P_{\rm sig}=-28$ dBm is chosen sufficiently low to avoid heating the on-chip attenuator. By sweeping the tone frequency, we extract the system noise [Fig. 3(b)], obtaining $29.3\pm 3.3$ photons at 5.5 GHz for an overall gain of $70.2\pm 0.5$ dB. We emphasize that the system noise found does not represent the intrinsic added noise of the HEMT ($\approx$ 5 photons at 5 K according to the datasheet), since the reference plane is taken at the output of the on-chip attenuator rather than at the HEMT input. This limitation can be removed by directly connecting the output of the on-chip attenuator to the isolator. For comparison, an independent calibration using the temperature sensor referenced to its input plane yields $14.7\pm 2.4$ photons (Supplementary Information S4).
@@ Acknowledgements.
This work was funded by the European Union’s HORIZON-RIA Programme under Grant No. 101135240 JOGATE. M.A.A. and S.S. acknowledge financial support from the European Union via Grant No. 101080167 ASPECTS. S.G. acknowledges financial support from the European Research Council (Grant No. 101041744 ESQuAT) and the Knut and Alice Wallenberg Foundation through the Wallenberg Centre for Quantum Technology (WACQT). The devices used in this work were fabricated in the Chalmers Myfab cleanroom facility. External interest disclosure: S.G. and L.A. are co-founders and equity holders in Sweden Quantum AB.
@@
The attenuation $A$ of the input line can be determined at the temperature sensor resonance frequency $\omega_{\rm ge}$, as detailed in ref. Scigliuzzo et al. (2020):

### v2

@@ Abstract
Figure 2: a) Sketch showing the frequency overlap between the thermal noise, the transmission band of the amplification chain, and the three output band-pass filters (1: 6.8-7.8 GHz, 2: 4.9-6.2 GHz; 3: 3.5-4.5 GHz) used to restrict the analog bandwidth in different power spectral density (PSD) measurements. For each filter, the signal tone to be calibrated, $\omega_{\rm sig}$, is swept in the rejection band of the filter (gray region), whereas noise is detected at $\omega_{\rm det}$ (black dot) in the transmission band (colored region). b) Power spectral densities measured at $\omega_{\rm det}=7$ GHz by Joule heating (purple) and a RF heating (blue) at $\omega_{\rm sig}=4$ GHz of the on-chip attenuator. The tone applied was rejected at room temperature with the output band-pass filter 1. The Joule and RF heating traces are separated by the attenuation AA of the input line. The tone heating was also acquired via the reference line without the on-chip attenuator (orange) to show the onset of heating of the bulkhead attenuators (diamond marker). Since their contribution is negligible (see main text), the best overlap (red) of the RF heating with the Joule heating giving the attenuation AA was fitted from $P_{\rm sig}=-22.5$ dBm (above the noise floor, dot marker) until an arbitrary power $P_{\rm sig}=-5$ dBm (cross marker). c) Attenuation profile obtained by comparing Joule and RF heating at different $\omega_{\rm sig}$. Three detection frequencies were considered and implemented with the set of output band-passes depicted in a). The red band in the background highlights the bandwidth of the amplification chain.
@@
where $N_{\rm sig}$ is the number of photons at frequency $\omega_{\rm sig}$ at the input of the amplification chain, SNR the signal to noise ratio and $P_{\rm input}=A_{\rm line}A_{\rm att}P_{\rm sig}$ the power at the input of the amplification chain. The tone power $P_{\rm sig}=-28$ dBm is chosen sufficiently low to avoid heating the on-chip attenuator. By sweeping the tone frequency, we extract the system noise [Fig. 3(b)], obtaining $23.6\pm 2.8$ photons at 5.5 GHz for an overall gain of $70.2\pm 0.5$ dB. We emphasize that the system noise found does not represent the intrinsic added noise of the HEMT ($\approx$ 5 photons at 5 K according to the datasheet), since the reference plane is taken at the output of the on-chip attenuator rather than at the HEMT input. This limitation can be removed by directly connecting the output of the on-chip attenuator to the isolator. For comparison, an independent calibration using the temperature sensor referenced to its input plane yields $14.2\pm 2.3$ photons (Supplementary Information S4).
@@ Acknowledgements.
This work was funded by the European Union’s HORIZON-RIA Programme under Grant No. 101135240 JOGATE. M.A.A. and $S.$ acknowledge financial support from the European Union via Grant No. 101080167 ASPECTS. S.G. acknowledges financial support from the European Research Council (Grant No. 101041744 ESQuAT) and the Knut and Alice Wallenberg Foundation through the Wallenberg Centre for Quantum Technology (WACQT). The devices used in this work were fabricated in the Chalmers Myfab cleanroom facility. External interest disclosure: S.G. and L.A. are co-founders and equity holders in Sweden Quantum AB.
@@
The attenuation AA of the input line can be determined at the temperature sensor resonance frequency $\omega_{\rm ge}$, as detailed in ref. Scigliuzzo et al. (2020):